# Robust Multimodal Learning with Missing Modalities via Parameter-Efficient Adaptation

### v4

@@ 5 Limitations and Future Directions
In this work, our main focus was to enhance missing modality robustness of existing multimodal models. Though our method can make existing models robust to different missing modality scenarios, it has certain limitations. First, we only considered missing modality during test time. However in real life scenarios, modalities can be missing in both train and test times. Second, we need to learn one set of adaptation parameters for every combination of missing modalities. While the number of adaptation parameters is small, the overall complexity can increase as the number of modality combinations increase. If we expect one modality to be missing at test time, adaptation parameter complexity would scale with the number of modalities. Third, we insert the learnable layers after each layer of the encoders and the fusion block. We did not try to optimize the number of parameters or find the optimal places to insert those learnable layers. Future study will explore these areas to further reduce the number of parameters, enhance the effectiveness and applicability of the approach in newer tasks and datasets.
@@ Acknowledgements
This work is supported in part by AFOSR award FA9550-21-1-0330 and NSF CAREER award CCF-2046293.
@@
S7 Why Dedicated Model Performs Better on MFNet Dataset?
@@
S8 Visualization of Predicted Segmentation Maps

### v6

@@ 4.3.3 Comparison with robust methods on NYUDv2 dataset
We observe that the CMNeXt model performs poorly when Depth is available and RGB is missing. This is due to its asymmetric architecture, which treats RGB as the primary modality and others as supplementary. As a result, performance drops significantly in the absence of RGB. However, the model overcomes this issue after adaptation and improves performance in all missing modality scenarios demonstrating the effectiveness of our adaptation approach.
@@ 5 Limitations and Future Directions
In this work, our main focus was to enhance missing modality robustness of existing multimodal models. Though our method can make existing models robust to different missing modality scenarios, it has certain limitations. First, we only considered missing modality during test time. However in real life scenarios, modalities can be missing in both train and test times. Second, our method learns one set of adaptation parameters for every combination of missing modalities. While the number of adaptation parameters is small, the overall parameter complexity will scale with the number of modality combinations. For $M$ modalities, we can have up to $2^{M}$ possible combinations, as each modality can either be available or missing. Our method will require $22^{M}-$ 2 sets of adaptation parameters to accommodate every possible combination of missing modalities (excluding two cases when all or none of the modalities are available). If we expect one modality out of $M$ to be missing at the test time, which is the case in most of the published work, our method will require $M$ sets of adaptation parameters. Third, we insert the learnable layers after each layer of the encoders and the fusion block. We did not try to optimize the number of parameters or find the optimal places to insert those learnable layers. Future study will explore these areas to further reduce the number of parameters, enhance the effectiveness and applicability of the approach in newer tasks and datasets.
@@ Acknowledgements
This work is supported in part by AFOSR award FA9550-21-1-0330 and NSF CAREER award CCF-2046293. This work used Indiana Jetstream2 through allocation CIS220128 from the ACCESS program supported by NSF grants 2138259, 2138286, 2138307, 2137603, and 2138296.
@@
Table S9: Missing modality performance comparison of base CMNeXt and Adapted CMNeXt model with Resnet-34 [72], Resnet-50 [72], and Swin-S [74] backbones on MFNet dataset.
Method Backbone Params RGB-Thermal RGB Thermal Missing Avg.
(M) mAcc %mIoU mAcc %mIoU mAcc %mIoU mAcc %mIoU
CMNeXt ResNet-34 50.72 50.56 45.61 23.95 16.98 19.25 16.94 21.60 16.96
Adapted CMNeXt ResNet-34 50.91 50.56 45.61 31.68 27.13 32.04 27.81 31.86 27.47
CMNeXt ResNet-50 54.68 50.69 46.05 12.00 11.18 19.45 18.38 15.73 14.78
Adapted CMNeXt ResNet-50 54.90 50.69 46.05 39.84 34.69 28.70 25.59 34.27 30.14
CMNeXt Swin-S 123.73 45.24 41.02 15.95 11.98 14.77 13.91 15.36 12.95
Adapted CMNeXt Swin-S 124.14 45.24 41.02 39.69 34.11 28.01 25.05 33.85 29.58
@@
S7 Effectiveness of our adaptation method on different backbones
To assess the effectiveness of our method across different backbones, we replaced the MiT-B4 backbone in the CMNeXt model with ResNet-34, ResNet-50, and Swin-S. We use the same experimental setup as the original CMNeXt model [61] while training the base models. In order to use the default window size and other configurations of Swin-S backbone, we resize the images to $448\times 448$ for training and testing. For ResNets, we use the default image resolution of $640\times 480$. Other hyperparameters are selected in the same manner as the original CMNeXt model. Then we adapt the pretrained base models for different missing modality scenarios using our approach. For adaptation, we use the same hyper-parameters as described in Section 4.2 and S2.
We summarize the test set performance of the pre-trained and adapted CMNeXt models on MFNet dataset with different backbones in Table S9. Our approach needs a small number of learnable parameters for each of the backbones. ResNet-34, ResNet-50, and Swin-S add only 0.37%, 0.40% and 0.33% additional learnable parameters, respectively. Adapted CMNeXt with ResNet-34 shows 10.26 and 10.51 points improvement in mean accuracy (mAcc) and %mIoU, respectively on average. Adapted CMNeXt with ResNet-50 provides an improvement of 18.54 and 15.36 points while the Adapted CMNeXt with Swin-S gains 18.49 and 16.63 points improvement in mAcc and %mIoU, respectively on average.
These results confirm that our adaptation approach generalizes across backbones, delivering significant performance gains with a small number of additional parameters.
S8 Why Dedicated Model Performs Better on MFNet Dataset?
@@
S9 Visualization of Predicted Segmentation Maps